Calculate the pH at 25oC of the following aqueous salt solutions and indicate if they are basic, acidic, or neutral.

a) 0.020 M KClO4(aq)

b)  0.50 M NaCN(aq)

c) 0.80 M NH4ClO4(aq)

Show all work please!
Answer #1

In solution, KClO4dissociates completely to K+and ClO4- ,the ClO4-causes hydrolysis forming HClO4 and OH-

ClO4- + H2O \rightleftharpoons HClO4 + OH-

therefore, we need the Kb of ClO4-

the Kb of HClO4 = 1 x 10-21

Kb = [HClO4] [OH-] / [ClO4-]

\Rightarrow1 x 10-21  = x2 / (0.02 - x)

\Rightarrow1 x 10-21  = x2 (0.02

\Rightarrow x2 = 2 x 10-23

\Rightarrow x = 4.4 x 10-12

[OH-] = x = 4.4 x 10-12 M

pOH = -log (4.4 x 10-12) = 11.36

pH = 14 - pOH = 14 - 11.36 = 2.64 (Acidic)

(b)

CN- + H2O \rightleftharpoons HCN + OH-

Kb = [HCN] [OH-] / [CN-] = 2.1 x 10^-5

Let
[HCH] = s
[OH-] = s
so, [CN-] = 0.5 - s

2.1 x 10-5 = (s)(s) / (0.5 - s)

Assume s << 0.011

2.1 x 10-5 = (s)(s) / (0.5)

s2 = 1.05 x 10-5

s = 3.24 x 10-3 M

[OH-] =s = 3.24 x 10-3 M

pOH = - log (3.24 x 10-3) = 2.49

pH + pOH = 14.00

pH = 14.00 - pOH = 14.00 - 2.49 = 11.51 (Basic)

(c)

NH4ClO4\rightleftharpoons NH4+ + ClO4-

Now in a water solution,
ClO4- + H2O \rightleftharpoons HClO4 + OH- (This doesn't happen because HClO4 is a strong acid. Therefore the only contribution to the pH is from the NH4+.

NH4+   \rightleftharpoons    NH3 + H+

IC 0.80 0 0

EC 0.80-x x x

Ka =  Kw/Kb = (1 x 10-14) / (1.8 x 10-5) = 5.5 x 10-10

Ka = [NH3] [H+] / [NH4+]

\Rightarrow5.5 x 10-10 = (x)(x)/(0.80-x)

\Rightarrow​5.5 x 10-10 = x2 /0.80 (Assume x << 0.80)

\Rightarrow​ x2 = 4.4 x 10-10

\Rightarrow​ x = 2.1 x 10-5

[H+] = x = 2.1 x 10-5

pH = -log(2.1 x 10-5) = 4.68 (Acidic)